OpenStudy (anonymous):

express the fraction 24/54 in lowest term

OpenStudy (anonymous):

Find common factors for each!

OpenStudy (anonymous):

\[\frac{ 24 }{ 54 }\div \frac{ 2 }{ 2 }=\frac{ 12 }{ 27 }\div \frac{ 3 }{ 3 }=\frac{ 4 }{ 9 }\] :o)

OpenStudy (anonymous):

Exactly. First you realize that 2 is a common factor of both 24 and 54 so you divide them both by 2. Then you realize that 3 is a common factor of 12 and 27 so you divide them both by 3. Then 4 and 9 have no common factor so you're done.

OpenStudy (anonymous):

Then divide by two to get: \[2\div4.5\] and you have your answer. gl ;D